OpenStudy (anonymous):

g(x) =12x^2 + 12x + 15 How do I find the solutions? I know it will have two real irrational solutions

OpenStudy (shamil98):

Find the gcf first. In this case it is 3. 3(4x^2 + 4x + 5)

OpenStudy (anonymous):

Oh, okay. I should have known that haha

OpenStudy (anonymous):

I know how to do this! it's just been a while, it's a review :/

OpenStudy (anonymous):

I still need help!

OpenStudy (shamil98):

You can use the quadratic formula for this.. x = -b +- √b^2 - 4ac / 2a

OpenStudy (shamil98):

I think doing that would be easier than factoring it out, since the solutions are complex.

OpenStudy (anonymous):

That formula doesn't look very familiar. Would I just plug in the numbers?

OpenStudy (shamil98):

ax^2 + bx + c = 0 those are the values of a b and c. have you taken algebra 1? you learn the quadratic formula then..

OpenStudy (anonymous):

I'm in Algebra 2! I'm really struggling though, and i have a hard time remembering formulas and steps. I'm completely familiar with ax^2 + bx + c = 0. And a few weeks ago we were learning about quadratic formulas, but what you showed me doesn't look familiar. Maybe we learned it a different way.

ganeshie8 (ganeshie8):

\(\large ax^2+bx+c=0\) quadratic formula to find solutions :- \(\large x~=~\frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

ganeshie8 (ganeshie8):

3(4x^2 + 4x + 5) = 0 \(\large 4x^2 + 4x + 5 = 0\) \(a = 4\) \(b = 4\) \(c = 5\) plug them above

OpenStudy (anonymous):

okay!! I think I understand that. But where does the distributive 3 go?

ganeshie8 (ganeshie8):

that goes away

ganeshie8 (ganeshie8):

\(3(4x^2 + 4x + 5) = 0\) \(3 \ne 0\), so, \(4x^2 + 4x + 5 = 0\)

ganeshie8 (ganeshie8):

look at below example :- \(100x = 0\) \(100 \ne 0\) so, \(x = 0\) where did the 100 go ?

OpenStudy (anonymous):

because 100 does not equal 0, it just simply disappears?

ganeshie8 (ganeshie8):

since 100 does not equal 0, the other guy must equal 0

OpenStudy (anonymous):

would the square root of -64 be 8i?

ganeshie8 (ganeshie8):

yup

OpenStudy (anonymous):

and -4 x 8i would be.. -32i?

ganeshie8 (ganeshie8):

careful, its a + or - sign in between not multiplication

OpenStudy (anonymous):

Oh! it would be addition, right?

ganeshie8 (ganeshie8):

-4 \(\pm\) 8i

OpenStudy (anonymous):

so -4 + 8i = 4i?

OpenStudy (anonymous):

and then 4i divided by 8?

ganeshie8 (ganeshie8):

noo

OpenStudy (anonymous):

:(

OpenStudy (shamil98):

-4 + 8i -4 - 8i those are your two answers ._.

OpenStudy (anonymous):

i'm hopeless.

ganeshie8 (ganeshie8):

divide by 2a

OpenStudy (anonymous):

i think i'm over thinking it.

OpenStudy (anonymous):

Well a is 4, right? and 2 times 4 is 8!

OpenStudy (shamil98):

-4 +- 8i / 2(4) = -4 +- 8i / 8

ganeshie8 (ganeshie8):

\(\large ax^2+bx+c=0\) quadratic formula to find solutions :- \(\large x~=~\frac{-b\pm\sqrt{b^2-4ac}}{2a}\) \(\large 4x^2 + 4x + 5\) \(\large x~=~\frac{-4\pm\sqrt{4^2-4 \times 4 \times 5}}{2 \times 4}\) \(\large x~=~\frac{-4\pm\sqrt{16-80}}{8}\) \(\large x~=~\frac{-4\pm\sqrt{-64}}{8}\) \(\large x~=~\frac{-4\pm 8i}{8}\) \(\large x~=~\frac{-1\pm 2i}{2}\)

OpenStudy (anonymous):

Ohhh, okay! So we're simplifying for the most part. Not solving.

ganeshie8 (ganeshie8):

dint get u

OpenStudy (shamil98):

thanks, ganshie, its too early in the morning for me to do quadratics and what not xD

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

I'm afraid to guess what the solutions would be.. it would be -1 +2i and -1 - 2i, right?

ganeshie8 (ganeshie8):

who ate the 2 in bottom ?

ganeshie8 (ganeshie8):

\(\large ax^2+bx+c=0\) quadratic formula to find solutions :- \(\large x~=~\frac{-b\pm\sqrt{b^2-4ac}}{2a}\) \(\large 4x^2 + 4x + 5\) \(\large x~=~\frac{-4\pm\sqrt{4^2-4 \times 4 \times 5}}{2 \times 4}\) \(\large x~=~\frac{-4\pm\sqrt{16-80}}{8}\) \(\large x~=~\frac{-4\pm\sqrt{-64}}{8}\) \(\large x~=~\frac{-4\pm 8i}{8}\) \(\large x~=~\frac{-1\pm 2i}{2}\) \(\large x~=~\frac{-1 + 2i}{2} ~~or ~~ x~=~\frac{-1 - 2i}{2} \)

OpenStudy (anonymous):

I've never seen it like that before. I would keep the denominator?

ganeshie8 (ganeshie8):

yup :)

OpenStudy (anonymous):

okay wow! that's what's completely throwing me off!

ganeshie8 (ganeshie8):

\(\large ax^2+bx+c=0\) quadratic formula to find solutions :- \(\large x~=~\frac{-b\pm\sqrt{b^2-4ac}}{2a}\) \(\large 4x^2 + 4x + 5\) \(\large x~=~\frac{-4\pm\sqrt{4^2-4 \times 4 \times 5}}{2 \times 4}\) \(\large x~=~\frac{-4\pm\sqrt{16-80}}{8}\) \(\large x~=~\frac{-4\pm\sqrt{-64}}{8}\) \(\large x~=~\frac{-4\pm 8i}{8}\) \(\large x~=~\frac{-1\pm 2i}{2}\) \(\large x~=~\frac{-1 + 2i}{2}~~ or~~ x~=~\frac{-1 - 2i}{2} \) \(\large x~=~\frac{-1}{2} + i ~~ or~~ x~=~\frac{-1}{2} - i \)

ganeshie8 (ganeshie8):

you may leave it like this also

OpenStudy (anonymous):

I like the other one better. :) but thank you so much!!! i can't believe you actually put up with me! haha. I really appreciate it.

ganeshie8 (ganeshie8):

lol you're good its fine :)